OpenStudy (anonymous):

whats the inverse of 3x-3?

OpenStudy (snuggielad):

What do you think it is? You have studied it so before we help yo tell us what you think

OpenStudy (anonymous):

okay this is what is its this 3+3y

OpenStudy (anonymous):

but i dont think its right

OpenStudy (anonymous):

umm.. y=3x-3, inverse, x=3y-3 3y=x+3 y=(1/3)x+1

OpenStudy (anonymous):

okay this is the equation that i have F(x)=3x-3 and i need the inverse of F-1(y)=???